Note vector notation: (vector 1;vector 2)


Select the correct answer:


k(k-1) • • • (k-r)(m+1)m• • • 3• 2

_________________________ =

p(p-1)• • • (p-m)(r+1)r• • • 3• 2• 


a) (k;r) / (p+1 ; m)


b) (p;r) / (k+1 ; m)


c) (k;r) / (p ; m + 1)


d) (k;r) / (m+1 ; p)


e) (k; r +1) / (p ; m)


f) (k; r + 1) / (p ; m+1)



1
Expert's answer
2021-05-28T10:19:41-0400

"\frac{k\left(k-1\right)\cdot \left(k-r\right)\left(m+1\right)m}{p\left(p-1\right)\cdot \left(p-m\right)\left(r+1\right)r}"


"=\frac{km\left(k-1\right)\left(-r+k\right)\left(m+1\right)}{r^2p^3+rp^3-r^2mp^2-rmp^2-r^2p^2-rp^2+r^2mp+rmp}"


"=\frac{k^3m^2+k^3m-k^2rm^2-k^2rm-k^2m^2-k^2m+krm^2+krm}{r^2p^3+rp^3-r^2mp^2-rmp^2-r^2p^2-rp^2+r^2mp+rmp}"


From the above expression, the answer is d) (k;r) / (m+1 ; p)